# Optimal Bias-Correction and Valid Inference in High-Dimensional Ridge Regression: A Closed-Form Solution

### v1

@@ 2.2 Ridge Regression
which simply replaces ${\mathbf{X}}^{\prime}{\mathbf{X}}$ by ${\mathbf{X}}^{\prime}{\mathbf{X}}+\lambda{\mathbf{I}}_{p}$ with a tuning parameter $\lambda>0$ in the least-squares estimator of (3). From a regression point of view, the ridge estimator can be obtained in the following way. Let ${\mathbf{y}}^{*}=({\mathbf{y}}^{\prime},{\bf 0})^{\prime}$ and ${\mathbf{X}}^{*}=({\mathbf{X}}^{\prime},\sqrt{\lambda}{\mathbf{I}}_{p})^{\prime}$ be the augmented data, the ridge estimator is a solution to the following optimization problem:
@@
which is the componentwise regression estimator if each covariate is standardized. Therefore, a large $\lambda$ would reduce the variance of the estimator, but the bias may increase as $\lambda$ grows.
@@
for any given $\lambda>0$ such that $({\mathbf{X}}{\mathbf{X}}^{\prime}+\lambda{\mathbf{I}}_{p})$ is invertible.
@@
The condition for the result in Theorem 1 to hold is the same as that for linear regression models. If the design matrix ${\mathbf{X}}$ is random with either independent and identically distributed ($i.i.d.$) or weakly dependent columns, we require $E(\mbox{\boldmath$\varepsilon$}|{\mathbf{X}})={\bf 0}$, and then, the bias of the ridge estimator conditioning on ${\mathbf{X}}$ and a given $\lambda>0$ is
@@
In the next section, we will tackle this issue by proposing an iterative procedure to reduce the bias of the ridge estimators.
@@ 2.3 Bias-Correction
The discussion in this section is divided into two key parts depending on whether ${\mathbf{X}}^{\prime}{\mathbf{X}}$ is invertible or singular. For simplicity, in line with the setting in Wang and Leng (2016), we assume that ${\mathbf{X}}^{\prime}{\mathbf{X}}$ is invertible when $p<n$ and singular when $p>n$ throughout this article [^1] An alternative framework is to consider the scenarios that $p/n\in(0,1)$ and $p/n\in(1,\infty)$, a common setting in random matrix theory, which is also helpful in establishing the asymptotic results if we further allow $n,p\rightarrow\infty$ later.. This assumption is well-justified as we consider a fixed design matrix ${\mathbf{X}}$ in this paper and ${\mathbf{X}}^{\prime}{\mathbf{X}}$ naturally maintains its invertibility when $p<n$. Furthermore, in extreme cases where highly correlated variables are present within ${\mathbf{X}}$ (in a random sense), we may implement specific transformations based on prior knowledge or statistical methods such as the hierarchical clustering or the $k$-means algorithm to mitigate these correlations prior to conducting ridge regression; see the discussion in Section 4.1.2 of Fan and Lv (2008). Consequently, we only focus on the bias-correction issue in this paper and rule out the case when some covariates in ${\mathbf{X}}$ are highly correlated.
Note that the ridge estimator in (5) can be written as
@@
It’s apparent that the $\ell_{2}$-norm of the bias term ${\mathbf{b}}_{\lambda,1}$ produced by $\widehat{\boldsymbol{\beta}}_{c,1}(\lambda)$ is smaller than that of the initial bias ${\mathbf{b}}_{\lambda,0}$ in Theorem 1 under mild conditions, implying that the bias ${\mathbf{b}}_{\lambda,0}$ has been partially corrected by $\widehat{\boldsymbol{\beta}}_{c,1}(\lambda)$. To see this, we conduct a singular-value decomposition on ${\mathbf{X}}$ or a spectral decomposition on ${\mathbf{X}}^{\prime}{\mathbf{X}}$, and the effectiveness of the bias-correction approach depends on two observations: (a) the eigenvalues of $\lambda({\mathbf{X}}^{\prime}{\mathbf{X}}+\lambda{\mathbf{I}}_{p})^{-1}$ are positive and strictly less than one; and (b) the eigenvalues of $\lambda^{2}({\mathbf{X}}^{\prime}{\mathbf{X}}+\lambda{\mathbf{I}}_{p})^{-2}$ becomes smaller in the de-biased estimator of (10) compared to those of $\lambda({\mathbf{X}}^{\prime}{\mathbf{X}}+\lambda{\mathbf{I}}_{p})^{-1}$ in ${\mathbf{b}}_{\lambda,0}$.
Following the first step, we replace the unknown vector $\boldsymbol{\beta}$ in (11) by the ridge estimator $\widehat{\boldsymbol{\beta}}(\lambda)$ again, leading us to construct a second-step de-biased estimator:
@@
where the eigenvalues of $\lambda^{3}({\mathbf{X}}^{\prime}{\mathbf{X}}+\lambda{\mathbf{I}}_{p})^{-3}$ are even smaller compared to those of $\lambda^{2}({\mathbf{X}}^{\prime}{\mathbf{X}}+\lambda{\mathbf{I}}_{p})^{-2}$ in ${\mathbf{b}}_{\lambda,1}$. Consequently, we can show that the $\ell_{2}$-norm of the bias ${\mathbf{b}}_{\lambda,2}$ is smaller than that of ${\mathbf{b}}_{\lambda,1}$ under the same framework. We continue this procedure and denote
@@
For $p<n$, ${\mathbf{X}}^{\prime}{\mathbf{X}}$ is invertible and the SVD of ${\mathbf{X}}$ is ${\mathbf{X}}={\mathbf{V}}_{1}{\mathbf{D}}_{1}{\mathbf{U}}_{1}^{\prime}$, where ${\mathbf{U}}_{1}^{\prime}{\mathbf{U}}_{1}={\mathbf{V}}_{1}^{\prime}{\mathbf{V}}_{1}={\mathbf{I}}_{p}$ and ${\mathbf{D}}_{1}=\mbox{diag}(d_{1},...,d_{p})$ with $d_{i}>0$ for $1\leq i\leq p$.
Assumption 1 is intuitive for a fixed design ${\mathbf{X}}$ with a large and fixed configuration of $(p,n)$. For example, the eigenvalues of ${\mathbf{X}}^{\prime}{\mathbf{X}}$ are of order $n$ if the entries of ${\mathbf{X}}$ are independent copies of a random variable with zero mean, unit variance, and finite fourth moment if $p/n\in(0,1)$; see the Bai-Yin’s law in Bai and Yin (1993). Consequently, the eigenvalues of $\lambda({\mathbf{X}}^{\prime}{\mathbf{X}}+\lambda{\mathbf{I}}_{p})^{-1}$ are strictly less than one, which ensures that our iterative procedure can substantially reduce the bias. In fact, we have the following theorem on the de-biased ridge estimator in (14).
@@
If $p<n$ and ${\mathbf{X}}^{\prime}{\mathbf{X}}$ is invertible. Under Assumption 1, the bias of the de-biased ridge estimator $\widehat{\boldsymbol{\beta}}_{c,k}(\lambda)$ defined in (14) is
@@
(ii) The requirement for the $\ell_{2}$-norm of $\boldsymbol{\beta}$ to be finite stems from the expectation of the response having a limited number of significant covariates in linear regression models. Without this restriction, as the number of covariates grows, each coefficient’s contribution might become sufficiently small, allowing the response’s variance to remain finite. It’s important to highlight that the condition for the number of iterations $k$ can still be met even if $\|\boldsymbol{\beta}\|_{2}\rightarrow\infty$ at a polynomial rate. This is because $(\frac{\lambda}{d_{j}^{2}+\lambda})^{k+1}$ decays exponentially, given that $|\frac{\lambda}{d_{j}^{2}+\lambda}|<1$ for $1\leq j\leq p$. Thus, the number of iterations $k$ can be selected to scale logarithmically with the dimension $p$.
(iii) As discussed in the Introduction section, the asymptotic results in Theorem 2 are established for any given $(p,n)$ as $k\rightarrow\infty$. This approach is also reasonable because it reflects common scenarios encountered in practical data analysis, where datasets often have fixed dimensions and sample sizes. As a matter of fact, if Assumption 1 holds for increasing $p$ and $n$ with $p<n$, the results remain applicable when $k\rightarrow\infty$ first, followed by $n,p\rightarrow\infty$. In addition, under the framework in Bai and Yin (1993) that $d_{i}^{2}\asymp n$, the asymptotic results hold simultaneously as $n,p,k\rightarrow\infty$ if $p/n\in(0,1)$ and the penalty parameter $\lambda\asymp n$ because $\max_{1\leq j\leq p}|\frac{\lambda}{d_{j}^{2}+\lambda}|<1$ still hold in this setting.
@@
To investigate the performance of the de-biased estimator $\widehat{\boldsymbol{\beta}}_{c,k}(\lambda)$ when $p>n$ under which ${\mathbf{X}}^{\prime}{\mathbf{X}}$ is a singular matrix, we first perform a singular-value-decomposition on ${\mathbf{X}}$. Suppose the true rank of ${\mathbf{X}}$ is $\text{rank}({\mathbf{X}})=p^{*}\leq\min(p,n)=n$, where we can simply set $p^{*}=n$ since we deal with a given and fixed configuration of $(p,n)$ without including highly correlated covariates. However, the results in Theorem 3 still hold for $p^{*}<n$. By abuse of notation, there exist semi-orthogonal matrices ${\mathbf{V}}_{1}\in R^{n\times p^{*}}$ and ${\mathbf{U}}_{1}\in R^{p\times p^{*}}$, and a diagonal matrix ${\mathbf{D}}_{1}=\mbox{diag}(d_{1},...,d_{p^{*}})$ with $d_{1}\geq...\geq d_{p^{*}}>0$ such that
@@
Rank(${\mathbf{X}}$)$=p^{*}\leq\min(p,n)=n$ if $p>n$, and the design matrix ${\mathbf{X}}$ has a SVD ${\mathbf{X}}={\mathbf{V}}_{1}{\mathbf{D}}_{1}{\mathbf{U}}_{1}^{\prime}$ such that ${\mathbf{X}}^{\prime}{\mathbf{X}}+\lambda{\mathbf{I}}_{p}={\mathbf{U}}{\mathbf{D}}{\mathbf{U}}^{\prime}$, where ${\mathbf{D}}$, ${\mathbf{D}}_{1}$, ${\mathbf{U}}$, ${\mathbf{U}}_{1}$, ${\mathbf{V}}_{1}$, and ${\mathbf{U}}_{2}$ are defined as those in (16) and (17).
@@
If $p>n$ and ${\mathbf{X}}^{\prime}{\mathbf{X}}$ is singular, but ${\mathbf{X}}^{\prime}{\mathbf{X}}+\lambda{\mathbf{I}}_{p}$ is invertible for a given $\lambda>0$. Under Assumption 2, the bias term ${\mathbf{b}}_{\lambda,k}$ of $\widehat{\boldsymbol{\beta}}_{c,k}(\lambda)$ is the same as (15) in Theorem 2. If the number of iterations $k$ satisfies $\max_{1\leq j\leq p^{*}}C_{n,p}(\frac{\lambda}{d_{j}^{2}+\lambda})^{k+1}\rightarrow 0$ where the constant $C_{n,p}$ is the same as that in Theorem 2, we have
@@
so long as $\max_{1\leq j\leq p^{*}}C_{n,p}(\frac{\lambda}{d_{j}^{2}+\lambda})^{k+1}\rightarrow 0$ and $d_{j}>0$ for $1\leq j\leq p^{*}$, where $p^{*}$ is the rank of ${\mathbf{X}}$.
(iv) It’s possible that the rank of $\mathbf{X}$ is $p^{*}<p$ in the case of $p<n$. In such situations, we can obtain similar results as those outlined in Theorem 3. The approach to handling this scenario is similar to that for $p>n$, and therefore, we omit further discussion on this and focus on the previously mentioned setting.
@@ 2.4 Ridge Screening
To further address the uncorrectable bias term identified in Theorem 3 and ensure valid statistical inferences when $p>n$, additional structures must be imposed on the parameters or covariates. Without these additional structures, as demonstrated in Theorem 4, the bias-correction becomes challenging. In the subsequent analysis, we will use $C$ or $c$ to denote a generic constant, the specific value of which may vary across different contexts.
In this section, we propose a ridge-screening approach to select the significant variables in Model (2). It is feasible for both scenarios when $p<n$ and $p>n$. Therefore, it can also be treated as a new variable selection approach especially useful in a high-dimensional setting, which is of independent interest to statisticians, econometricians, and data scientists. We only focus on the scenario when $p>n$ and ${\mathbf{X}}{\mathbf{X}}^{\prime}$ is singular in such a situation. Note that the reason for the remaining bias term in Theorem 3 that cannot be corrected is that some projection directions of the coefficients $\boldsymbol{\beta}$ are not captured in Model (2). For example, this is the case when some parameters in $\boldsymbol{\beta}$ are redundant if certain covariates in ${\mathbf{X}}$ are strongly correlated. Therefore, it is reasonable to make an assumption that some covariates in ${\mathbf{X}}$ are not useful in the linear regression (2), and hence, they can be dropped first before establishing valid ridge estimators. To embrace the sparsity assumption in high-dimensional data analysis, we assume the true parameter vector $\boldsymbol{\beta}$ belongs to the following submodel class
$\mathcal{M}_{0}=\{1\leq i\leq p,|\beta_{i}|\,\,\text{is not zero}\},$ (18)
where we assume its cardinality $|\mathcal{M}_{0}|=s^{*}<\min(p,n)$. It is obvious that we require $s^{*}<p^{*}$, which is the rank of ${\mathbf{X}}$ defined in Assumption 2. This cardinality assumption is natural and it reflects the idea that many coefficient parameters are relatively small among the $p$-dimensional vector $\boldsymbol{\beta}$. We treat them as zero elements only for ease of exploitation. According to the results in Theorem 2 and Theorem 3, we can see that
@@
where $\widehat{\boldsymbol{\beta}}_{c,k}(\lambda^{*})=(\widehat{\beta}_{c,k,1}(\lambda^{*}),...,\widehat{\beta}_{c,k,p}(\lambda^{*}))^{\prime}$ and we use a different penalty $\lambda^{*}>0$ to distinguish it from the one utilized in the subsequent step. The ranking method to derive the submodel in (20) is similar to the approach presented in Section 2.2 of Fan and Lv (2008). However, the method in Fan and Lv (2008) is based on marginal correlations between the response and the features, while (20) is a utilization of a de-biased estimator, which is more proximate to the true parameter than the conventional ridge estimator in Eq. (5) of Fan and Lv (2008). Furthermore, our methodology is not constrained by a specific choice of $\lambda^{*}$, and it can be chosen by an information criterion or a cross-validation method as discussed at the end of Section 2.4.
In practice, we may select $n^{*}<\min(p,n)=n$ in (20) through cross-validation along with $\lambda^{*}$ because the actual design matrix with $n^{*}$ columns of covariates is no longer singular in such cases. Additionally, we can show that the model with $n^{*}$ covariates asymptotically encompasses the submodel $\mathcal{M}_{0}$ in (18) under mild conditions, i.e., the true submodel $\mathcal{M}_{0}$ in (18) is nested within the one with $n^{*}$ covariates.
Now, we restrict the design matrix to the submodel class $\mathcal{M}_{k}(\lambda^{*})$ and denote the restricted design as ${\mathbf{X}}_{\mathcal{M}_{k}}\in R^{n\times n^{*}}$, and the new ridge estimator is
@@
To establish the asymptotic properties of the RS method and the de-biased estimator for the restricted one in (22), we make a few intuitive assumptions first.
@@
For any submatrix ${\mathbf{X}}_{\mathcal{M}_{k}}$ of ${\mathbf{X}}$ with dimension $n^{*}<n$, all the eigenvalues of ${\mathbf{X}}_{\mathcal{M}_{k}}^{\prime}{\mathbf{X}}_{\mathcal{M}_{k}}$, denoted as $d_{\mathcal{M}_{k},j}^{2}$ for $1\leq j\leq n^{*}$, are of order $n$.
@@
For $i\in\mathcal{M}_{0}$ defined in (18), $\min_{i\in\mathcal{M}_{0}}|\beta_{i}|\geq C_{1}n^{-\tau}$ for some $0<\tau<n$, and the magnitude of the $i$-th projected coordinate $|({\mathbf{U}}_{2}{\mathbf{U}}_{2}^{\prime}\boldsymbol{\beta})_{i}|\leq C_{2}|\beta_{i}|$ for $C_{2}<1$, and $\|\boldsymbol{\beta}\|_{2}^{2}\leq C_{3}s^{*}$, where $s^{*}$ is the number of nonzero elements in $\boldsymbol{\beta}$.

### v2

@@ 2.2 Ridge Regression
which simply replaces ${\mathbf{X}}^{\prime}{\mathbf{X}}$ by ${\mathbf{X}}^{\prime}{\mathbf{X}}+\lambda{\mathbf{I}}_{p}$ with a tuning parameter $\lambda>0$ in the least-squares estimator of (3). From a regression point of view, the ridge estimator can be obtained in the following way. Let ${\mathbf{y}}^{*}=({\mathbf{y}}^{\prime},{\bf 0})^{\prime}$ and ${\mathbf{X}}^{*}=({\mathbf{X}}^{\prime},\sqrt{\lambda}{\mathbf{I}}_{p})^{\prime}$ be the augmented data, the ridge estimator is a solution to the following optimisation problem:
@@
which is the componentwise regression estimator if each covariate is standardised. Therefore, a large $\lambda$ would reduce the variance of the estimator, but may increase the bias. In fact, the bias increases as $\lambda$ grows.
@@
for any given $\lambda>0$ such that $({\mathbf{X}}^{\prime}{\mathbf{X}}+\lambda{\mathbf{I}}_{p})$ is invertible.
@@
The condition for the result of Theorem 1 to hold is the same as that for the linear regression model. If the design matrix ${\mathbf{X}}$ is random with either independent and identically distributed ($i.i.d.$) or weakly dependent columns, we require $E(\mbox{\boldmath$\varepsilon$}|{\mathbf{X}})={\bf 0}$, and then, the bias of the ridge estimator conditioning on ${\mathbf{X}}$ and a given $\lambda>0$ is
@@
In the next section, we will tackle this issue by proposing an iterative procedure to reduce the bias of the ridge estimator.
@@ 2.3 Bias-Correction
We divide the discussion into two key parts depending on whether ${\mathbf{X}}^{\prime}{\mathbf{X}}$ is invertible or singular. For simplicity and in line with the setting in Wang and Leng (2016), we assume that ${\mathbf{X}}^{\prime}{\mathbf{X}}$ is invertible when $p<n$ and singular when $p>n$ throughout this article. An alternative framework is to consider the scenarios that $p/n\in(0,1)$ and $p/n\in(1,\infty)$, a common setting in random matrix theory, which is also helpful in establishing the asymptotic results if we further allow $n,p\rightarrow\infty$ later. This assumption is well-justified as we consider a fixed design matrix ${\mathbf{X}}$ in this paper and ${\mathbf{X}}^{\prime}{\mathbf{X}}$ naturally maintains its invertibility when $p<n$. Furthermore, in extreme cases where highly correlated variables are present within ${\mathbf{X}}$ (in a random sense), we may implement specific transformations based on prior knowledge or statistical methods such as the hierarchical clustering or the $k$-means algorithm to mitigate these correlations prior to conducting ridge regression; see the discussion in Section 4.1.2 of Fan and Lv (2008). Consequently, we only focus on the bias-correction issue in this paper and rule out the case when some covariates in ${\mathbf{X}}$ are highly correlated.
The ridge estimator in (5) can be written as
@@
It is then apparent that the $\ell_{2}$-norm of the bias term ${\mathbf{b}}_{\lambda,1}$ produced by $\widehat{\boldsymbol{\beta}}_{c,1}(\lambda)$ is smaller than that of the initial bias ${\mathbf{b}}_{\lambda,0}$ in Theorem 1 under some mild conditions, implying that the bias ${\mathbf{b}}_{\lambda,0}$ has been partially corrected by $\widehat{\boldsymbol{\beta}}_{c,1}(\lambda)$. To see this, we conduct a singular-value decomposition on ${\mathbf{X}}$ or a spectral decomposition on ${\mathbf{X}}^{\prime}{\mathbf{X}}$, and the effectiveness of the bias-correction approach depends on two observations: (a) the eigenvalues of $\lambda({\mathbf{X}}^{\prime}{\mathbf{X}}+\lambda{\mathbf{I}}_{p})^{-1}$ are positive and strictly less than one; and (b) the eigenvalues of $\lambda^{2}({\mathbf{X}}^{\prime}{\mathbf{X}}+\lambda{\mathbf{I}}_{p})^{-2}$ becomes smaller in the de-biased estimator of (10) compared to those of $\lambda({\mathbf{X}}^{\prime}{\mathbf{X}}+\lambda{\mathbf{I}}_{p})^{-1}$ in ${\mathbf{b}}_{\lambda,0}$.
Following the first correction, we replace the unknown vector $\boldsymbol{\beta}$ in (11) by the ridge estimator $\widehat{\boldsymbol{\beta}}(\lambda)$ again to obtain a second-step de-biased estimator:
@@
where the eigenvalues of $\lambda^{3}({\mathbf{X}}^{\prime}{\mathbf{X}}+\lambda{\mathbf{I}}_{p})^{-3}$ are even smaller compared to those of $\lambda^{2}({\mathbf{X}}^{\prime}{\mathbf{X}}+\lambda{\mathbf{I}}_{p})^{-2}$ in ${\mathbf{b}}_{\lambda,1}$. Consequently, we can show that the $\ell_{2}$-norm of the bias ${\mathbf{b}}_{\lambda,2}$ is smaller than that of ${\mathbf{b}}_{\lambda,1}$ under the same framework as before. We repeat this procedure and denote
@@
For $p<n$, ${\mathbf{X}}^{\prime}{\mathbf{X}}$ is invertible and the SVD of ${\mathbf{X}}$ is ${\mathbf{X}}={\mathbf{V}}_{1}{\mathbf{D}}_{1}{\mathbf{U}}_{1}^{\prime}$, where ${\mathbf{U}}_{1}^{\prime}{\mathbf{U}}_{1}={\mathbf{V}}_{1}^{\prime}{\mathbf{V}}_{1}={\mathbf{I}}_{p}$ and ${\mathbf{D}}_{1}=\mbox{diag}(d_{1},...,d_{p})$ with $d_{i}>0$, for $1\leq i\leq p$.
Assumption 1 is intuitive for a fixed design ${\mathbf{X}}$ with a large and fixed configuration of $(p,n)$. For example, the eigenvalues of ${\mathbf{X}}^{\prime}{\mathbf{X}}$ are of order $n$ if the entries of ${\mathbf{X}}$ are independent copies of a random variable with zero mean, unit variance, and finite fourth moment if $p/n\in(0,1)$; see the Bai-Yin’s law in Bai and Yin (1993). Consequently, the eigenvalues of $\lambda({\mathbf{X}}^{\prime}{\mathbf{X}}+\lambda{\mathbf{I}}_{p})^{-1}$ are strictly less than one, which ensures that our iterative procedure can substantially reduce the bias. In fact, we have the following theorem for the de-biased ridge estimator in (14).
@@
If $p<n$ and ${\mathbf{X}}^{\prime}{\mathbf{X}}$ is invertible, then under Assumption 1, the bias of the de-biased ridge estimator $\widehat{\boldsymbol{\beta}}_{c,k}(\lambda)$ defined in (14) is
@@
(ii) The requirement for the $\ell_{2}$-norm of $\boldsymbol{\beta}$ to be finite stems from the expectation of the response having a limited number of significant covariates in linear regression models. Without this restriction, as the number of covariates grows, each coefficient’s contribution might become sufficiently small, allowing the response’s variance to remain finite. It’s important to highlight that the condition for the number of iterations $k$ can still be met even if $\|\boldsymbol{\beta}\|_{2}\rightarrow\infty$ at a polynomial rate. This is because $(\frac{\lambda}{d_{j}^{2}+\lambda})^{k+1}$ decays exponentially, given that $|\frac{\lambda}{d_{j}^{2}+\lambda}|<1$, for $1\leq j\leq p$. Thus, the number of iterations $k$ can be selected to scale logarithmically with the dimension $p$.
(iii) As discussed in the Introduction section, the asymptotic results in Theorem 2 are established for any given $(p,n)$ as $k\rightarrow\infty$. This approach is also reasonable because it reflects common scenarios encountered in practical data analysis, where datasets often have fixed dimensions and sample sizes. As a matter of fact, if Assumption 1 holds for increasing $p$ and $n$ with $p<n$, the results remain applicable when $k\rightarrow\infty$ first, followed by $n,p\rightarrow\infty$. In addition, under the framework in Bai and Yin (1993) that $d_{i}^{2}\asymp n$, the asymptotic results hold simultaneously as $n,p,k\rightarrow\infty$ if $p/n\in(0,1)$ and the penalty parameter $\lambda\asymp n$ because $\max_{1\leq j\leq p}|\frac{\lambda}{d_{j}^{2}+\lambda}|<1$ still holds in this setting.
@@
To investigate the performance of the de-biased estimator $\widehat{\boldsymbol{\beta}}_{c,k}(\lambda)$ when $p>n$ under which ${\mathbf{X}}^{\prime}{\mathbf{X}}$ is singular, we first perform a singular-value-decomposition on ${\mathbf{X}}$. Suppose the true rank of ${\mathbf{X}}$ is $\text{rank}({\mathbf{X}})=p^{*}\leq\min(p,n)=n$, where we can simply set $p^{*}=n$ since we deal with a given and fixed configuration of $(p,n)$ without including highly correlated covariates. However, the results in Theorem 3 still hold for $p^{*}<n$. By abuse of notation, there exist semi-orthogonal matrices ${\mathbf{V}}_{1}\in R^{n\times p^{*}}$ and ${\mathbf{U}}_{1}\in R^{p\times p^{*}}$, and a diagonal matrix ${\mathbf{D}}_{1}=\mbox{diag}(d_{1},...,d_{p^{*}})$ with $d_{1}\geq...\geq d_{p^{*}}>0$ such that
@@
Rank(${\mathbf{X}}$)$=p^{*}\leq\min(p,n)=n$ if $p>n$, and the design matrix ${\mathbf{X}}$ has a SVD ${\mathbf{X}}={\mathbf{V}}_{1}{\mathbf{D}}_{1}{\mathbf{U}}_{1}^{\prime}$ such that ${\mathbf{X}}^{\prime}{\mathbf{X}}+\lambda{\mathbf{I}}_{p}={\mathbf{U}}{\mathbf{D}}{\mathbf{U}}^{\prime}$, where ${\mathbf{D}}$, ${\mathbf{D}}_{1}$, ${\mathbf{U}}$, ${\mathbf{U}}_{1}$, ${\mathbf{V}}_{1}$, and ${\mathbf{U}}_{2}$ are defined in (16) and (17).
@@
If $p>n$ and ${\mathbf{X}}^{\prime}{\mathbf{X}}$ is singular, but ${\mathbf{X}}^{\prime}{\mathbf{X}}+\lambda{\mathbf{I}}_{p}$ is invertible for a given $\lambda>0$. Under Assumption 2, the bias term ${\mathbf{b}}_{\lambda,k}$ of $\widehat{\boldsymbol{\beta}}_{c,k}(\lambda)$ is the same as (15) in Theorem 2. If the number of iterations $k$ satisfies $\max_{1\leq j\leq p^{*}}C_{n,p}(\frac{\lambda}{d_{j}^{2}+\lambda})^{k+1}\rightarrow 0$, where the constant $C_{n,p}$ is the same as that in Theorem 2, we have
@@
so long as $\max_{1\leq j\leq p^{*}}C_{n,p}(\frac{\lambda}{d_{j}^{2}+\lambda})^{k+1}\rightarrow 0$ and $d_{j}>0$, for $1\leq j\leq p^{*}$, where $p^{*}$ is the rank of ${\mathbf{X}}$.
(iv) It is possible that the rank of $\mathbf{X}$ is $p^{*}<p$ in the case of $p<n$. In such situations, we can obtain similar results as those outlined in Theorem 3. The approach to handling this scenario is similar to that for $p>n$, and therefore, we omit further discussion on this issue and focus on the previously mentioned setting.
@@ 2.4 Ridge Screening
To address further the uncorrectable bias identified in Theorem 3 and to ensure valid statistical inferences when $p>n$, we impose some additional structures on the parameters or covariates. Without these additional structures, as demonstrated in Theorem 4, the bias-correction becomes challenging. In the subsequent analysis, we use $C$ or $c$ to denote a generic constant, the specific value of which may vary across different contexts.
In this section, we propose a ridge-screening approach to select the significant variables in Model (2). It is feasible for both scenarios when $p<n$ and $p>n$. Therefore, it can also be treated as a new variable selection approach especially useful in a high-dimensional setting, which is of independent interest to statisticians, econometricians, and data scientists. We only focus on the scenario when $p>n$ and ${\mathbf{X}}{\mathbf{X}}^{\prime}$ is singular. Note that the reason for the remaining bias term in Theorem 3 that cannot be corrected is that some projection directions of the coefficients $\boldsymbol{\beta}$ are not captured in Model (2). For example, this is the case when some parameters in $\boldsymbol{\beta}$ are redundant if certain covariates in ${\mathbf{X}}$ are strongly correlated. Therefore, it is reasonable to assume that some covariates in ${\mathbf{X}}$ are not useful in the linear regression (2), and hence, they can be dropped before establishing valid ridge estimators. To embrace the sparsity assumption in high-dimensional data analysis, we assume the true parameter vector $\boldsymbol{\beta}$ belongs to the following submodel class
$\mathcal{M}_{0}=\{1\leq i\leq p,|\beta_{i}|>\kappa\},$ (18)
with cardinality $|\mathcal{M}_{0}|=s^{*}<\min(p,n)$, where $\kappa$ is a small positive number. It is obvious that we require $s^{*}<p^{*}$, which is the rank of ${\mathbf{X}}$ defined in Assumption 2. This cardinality assumption is natural and it reflects the idea that many coefficient parameters are relatively small among the $p$-dimensional vector $\boldsymbol{\beta}$. We treat them as zero elements only for ease of exploitation. According to the results in Theorem 2 and Theorem 3, we have that
@@
where $\widehat{\boldsymbol{\beta}}_{c,k}(\lambda^{*})=(\widehat{\beta}_{c,k,1}(\lambda^{*}),...,\widehat{\beta}_{c,k,p}(\lambda^{*}))^{\prime}$ and we use a different penalty $\lambda^{*}>0$ to distinguish it from the one utilised in the subsequent step. We clarify the notation used here. For $p>n$, $s^{*}$ is the number of nonzero elements in $\boldsymbol{\beta}$, $p^{*}$ denotes the rank of ${\mathbf{X}}$, and $n^{*}$ is the number of covariates selected by the RS method. Obviously, we require that $s^{*}<n^{*}\leq p^{*}\leq\min(p,n)=n$ for $p>n$. We can also simply set $p^{*}=n$ since we deal with a fixed design.
The ranking method to derive the submodel in (20) is similar to the approach presented in Section 2.2 of Fan and Lv (2008). However, the method in Fan and Lv (2008) is based on marginal correlations between the response and the features, while (20) is a utilisation of a de-biased estimator, which is more proximate to the true parameter than the conventional ridge estimator in Eq. (5) of Fan and Lv (2008). Furthermore, our methodology is not constrained by a specific choice of $\lambda^{*}$, and the optimal one can be chosen by an information criterion or a cross-validation method as discussed at the end of Section 2.4.
In practice, we may select $n^{*}<\min(p,n)=n$ in (20) through cross-validation along with $\lambda^{*}$ because the actual design matrix with $n^{*}$ columns of covariates is no longer singular in such cases. Additionally, we can show that the model with $n^{*}$ covariates asymptotically encompasses the submodel $\mathcal{M}_{0}$ in (18) under some mild conditions, i.e., the true submodel $\mathcal{M}_{0}$ in (18) is nested within the one with $n^{*}$ covariates.
Next, we restrict the design matrix to the submodel class $\mathcal{M}_{k}(\lambda^{*})$ and denote the restricted design as ${\mathbf{X}}_{\mathcal{M}_{k}}\in R^{n\times n^{*}}$. The new ridge estimator becomes
@@
To establish the asymptotic properties of the RS method and the de-biased estimator for the restricted one in (22), we make a few intuitive assumptions below.
@@
For any submatrix ${\mathbf{X}}_{\mathcal{M}_{k}}$ of ${\mathbf{X}}$ with dimension $n^{*}<n$, all the eigenvalues of ${\mathbf{X}}_{\mathcal{M}_{k}}^{\prime}{\mathbf{X}}_{\mathcal{M}_{k}}$, denoted as $d_{\mathcal{M}_{k},j}^{2}$, for $1\leq j\leq n^{*}$, are of order $n$.
@@
For $i\in\mathcal{M}_{0}$ defined in (18), $\min_{i\in\mathcal{M}_{0}}|\beta_{i}|\geq\kappa\geq C_{1}n^{-\tau}$ for some $20<\tau<1/$ 2, and the magnitude of the $i$-th projected coordinate $|({\mathbf{U}}_{2}{\mathbf{U}}_{2}^{\prime}\boldsymbol{\beta})_{i}|\leq C_{2}|\beta_{i}|$ for $C_{2}<1$, and $\|\boldsymbol{\beta}\|_{2}^{2}\leq C_{3}s^{*}$, where $s^{*}$ is the number of nonzero elements in $\boldsymbol{\beta}$.